OpenStudy (anonymous):

log[4](xy)^3 - log[4](xy) =3log[4]x + 3log[4]y - (log[4]x + log[4] y) =2log[4]x + 2log[4]y =2log[4](xy) =log[4](xy)^2 is this correct?

OpenStudy (callisto):

Yes, but do you get the same answer in 2 steps.. \[\log_{4} (xy)^3 -\log_{4} (xy)\]\[=\log_{4}[ (xy)^3 /(xy)]\]\[=\log_{4} (xy)^2 \]

OpenStudy (callisto):

Note that \[loga - logb = \log (a/b)\]

OpenStudy (anonymous):

Thankx ;)

OpenStudy (callisto):

welcome